# Establishing Task Scaling Laws via Compute-Efficient Model Ladders

### v1

@@ B.2 Predicting task accuracy under the MC format
In this paper we have been focusing on the ranked classification (RC) format of tasks. Here we explore if we can make predictions when problems are written in the multiple-choice (MC) format. Unfortunately, all ladder models are small and show random performance in MC on all tasks (Gu et al., 2024), and thus it is not practical to fit the mapping from task loss to MC accuracy in step 2 with data points from the ladder models. Instead, we use data points from early intermediate checkpoints of the target models.
@@ Appendix C Additional Details on Design Choices
C.1 Task loss prediction: The impact of using compute-FLOPs instead of
Figure 11 shows the function fitting and prediction in step 1, using compute-FLOPs as the input feature.
C.2 Task cross-entropy as intermediate feature
Our intermediate feature in step 1 is the conditional loss of the correct answer of a task, from which we predict the accuracy in step 2. However, the accuracy is computed from the losses of both correct and incorrect continuations, as seen in Equation 5.
We commonly maximize the accuracy by minimizing the task cross-entropy as a surrogate loss, in which the terms are used as logits:
We refer to this as TaskCE to distinguish it from the language modeling cross-entropy over tokens. We investigate this cross entropy as an alternative intermediate feature, which takes into account the incorrect answer choices as an additional log partition term.
Step 2 scaling with Task Cross-Entropy.
@@ Results.
While the average fitting and extrapolation errors in step 1 of Figure 12 are moderate and comparable to Figure 3. However, we notice that the fitting errors for task cross-entropy tend to be more skewed, i.e., the fitted Chinchilla curves tend to underestimate the performance of the 190M-1xC model and overstimate the 190M-10xC, while underestimating the 1B-1xC and overestimating the 1B-10xC. As these errors appear more systematic, it raises questions whether this parametric form is a good fit over much larger scales. Notably, it results in substantial errors in the task cross-entropy of MMLU and ARC-Challenge (4.2% - 4.8%).
@@
Finally, we note that the values for task cross-entropy fall into a narrow range. This means that the predictions in step 2 are highly sensitive to small changes in the task cross-entropy, e.g., a difference in task cross-entropy of less than nats in PIQA can make the difference between random task accuracy (at ) and 83% accuracy (at ). As a consequence, small relative errors in step 1 (0.6% for the 13B model on PIQA) are amplified in Step 2 and result in larger overall errors in accuracy (2.9% for PIQA), despite a good fit in step 2 (0.9% prediction error for the 13B accuracy). Similarly, a step-1 prediction error of 4.7% for the 13B model on ARC-Challenge results in an overall error of 19.5%.
C.3 General language modeling loss as intermediate feature
Figure 15 shows the function fitting and predictions using the C4 language modeling loss as intermediate feature (§6.3).
@@ C.3 General language modeling loss as intermediate feature
In step 1, the fitted function underestimates the C4 loss by 2.0% for 7B-4T and 3.8% for 13B-5T. This error is higher than that on task loss prediction for 4 out of 8 tasks. The step 2 error is also higher than using task loss on 4 out of 8 tasks. When chaining the two steps to predict task accuracy, using C4 loss as intermediate feature resulted in higher error on 5 tasks – MMLU, HellaSwag, CommonsenseQA, Social IQA, and OpenBookQA.
C.4 Combining step 1 and 2 into a single step
@@
The prediction error is higher on 4 out of 8 tasks – MMLU, ARC-Challenge, CommonsenseQA, and Social IQA. In particular, on CommonsenseQA the fitted functions does not vary with respect to model size, indicating a degenerated function fitting. We conclude that the single-step approach is not as robust as the two-step approach.

### v2

@@ B.2 Predicting task accuracy under the MC format
In this paper we have been focusing on the ranked classification (RC) format of tasks. Here, we explore if we can make predictions when problems are written in the multiple-choice (MC) format. This is a challenging problem, as small models often exhibit random performance on standard downstream tasks (e.g., MMLU) until a certain size threshold (Hu et al., 2023). Thus, it is not practical to fit the mapping from task loss to MC accuracy in step 2 with data points from the ladder models. Instead, we use data points from early intermediate checkpoints of the target models.
@@
B.3 Compute requirements for predicting each task
We consider the impact of the scale of the model ladder on the prediction for the target model. In general, the prediction is harder if the difference of scale between the ladder and the target models is larger. We explore this trade-off using the 7B-4T model as the target.
In Figure 8, we progressively increase the compute-FLOPs used for function fitting. The prediction error is significantly worse with fewer FLOPs. On MMLU, our full ladder (3.2% compute of the target model) gets 1.3% error. Reducing the number of FLOPs to 0.1% of the target model compute increases the error to 12% which is an order of magnitude higher. This is more observable in certain tasks. Interestingly, we see a slight improvement in errors with less compute in ARC-C and ARC-E, potentially due to the variance as seen in §5.
Figure 10 similarly shows the impact of including models with longer training regimes. Interestingly, we do not see a significant reduction in the prediction error as we include the ladder models trained to higher Chinchilla multipliers444The target model 7B-4T is trained to approximately 28xC.. There is a slight downward trend (noticeable in MMLU), but this dimension has more flexibility.
@@ Appendix C Additional Details on Design Choices
C.1 Compute-FLOPs instead of for task loss prediction
We test if compute-FLOPs can be used directly for task loss prediction instead of in the over-trained regime with the ladder models. We fit a similar power function
where are parameters to fit.
Overall, compute-FLOPs are not expressive enough to distinguish between compute-optimal and overtrained models with equal FLOPs but different losses (Hoffmann et al., 2022). Thus, we do not use it in our main method.
C.2 Task cross-entropy as the intermediate feature
Task loss only considers the correct choice of each problem. However, task RC accuracy is determined by the losses of both correct and incorrect choices, which is a major challenge for predicting downstream performance (Schaeffer et al., 2024):
where is the loss over the ’th answer option of the ’th example and is the label of the correct answer 555Loss terms may include some normalization, e.g., by the number of characters in each answer or the unconditional answer probability (Gu et al., 2024)..
To account for incorrect answers, we define an alternative intermediate feature. We commonly maximize the accuracy by minimizing the task cross-entropy as a surrogate loss, in which the terms are used as logits:
We refer to this as TaskCE to distinguish it from the language modeling cross-entropy over tokens.
@@ C.2 Task cross-entropy as the intermediate feature
Step 2 scaling with Task Cross-Entropy.
@@ Results.
The average fitting and extrapolation errors in step 1 of Figure 12 are moderate and comparable to Figure 3. However, we notice that the fitting errors for TaskCE tend to be more skewed, i.e., the fitted Chinchilla curves tend to underestimate the performance of the 190M-1xC model and overstimate the 190M-10xC, while underestimating the 1B-1xC and overestimating the 1B-10xC. As these errors appear more systematic, it raises questions whether this parametric form is a good fit over much larger scales. Notably, it results in substantial errors in the task cross-entropy of MMLU and ARC-Challenge (4.2% - 4.8%).
@@
Finally, we note that the values for TaskCE fall into a narrow range. This means that the predictions in step 2 are highly sensitive to small changes in the task cross-entropy, e.g., a difference in task cross-entropy of less than nats in PIQA can make the difference between random task accuracy (at ) and 83% accuracy (at ). As a consequence, small relative errors in step 1 (0.6% for the 13B model on PIQA) are amplified in Step 2 and result in larger overall errors in accuracy (2.9% for PIQA), despite a good fit in step 2 (0.9% prediction error for the 13B accuracy). Similarly, a step-1 prediction error of 4.7% for the 13B model on ARC-Challenge results in an overall error of 19.5%.
C.3 General language modeling loss as the intermediate feature
Language modeling loss (LM loss) on held-out sets has been shown to follow the power law (Hoffmann et al., 2022). Here, we consider if we can map it to task performance.
We experiment with using the LM loss on the C4-en validation set (Raffel et al., 2019) as the intermediate feature.
Figure 15 shows the function fitting and predictions.
@@ C.3 General language modeling loss as the intermediate feature
In step 1, the fitted function underestimates the C4 loss by 2.0% for 7B-4T and 3.8% for 13B-5T. This error is higher than that on task loss prediction for 4 out of 8 tasks. The step 2 error is also higher than using task loss on 4 out of 8 tasks. When chaining the two steps to predict task accuracy, using C4 loss as intermediate feature resulted in higher error on 5 tasks – MMLU, HellaSwag, CommonsenseQA, Social IQA, and OpenBookQA. Prediction error is lower on ARC-Challenge and ARC-Easy. We conclude that using C4 loss can benefit certain tasks where task loss does not work well.
C.4 Comparison of three intermediate features
C.5 Combining step 1 and 2 into a single step
Here, we try to directly predict task accuracy from the training scale in one step, by combining Equation 1 and Equation 2 into a single parameterized function, and merging parameters and into , so that we reduce to 7 free parameters:
With this, we remove the a priori definition of a specific intermediate feature (i.e., the expression no longer carries a specific meaning), while preserving the representation power of the function. This function can be harder to fit, since it has more free parameters, is not convex, and we cannot use any data points collected from intermediate checkpoints as we did in step 2.
@@ C.5 Combining step 1 and 2 into a single step
We fit this function with data points from the final checkpoints of the ladder models, using the same optimization method as in step 1 (Figure 17). The prediction error is higher on 4 out of 8 tasks – MMLU, ARC-Challenge, CommonsenseQA, and Social IQA. In particular, on CommonsenseQA the fitted functions does not vary with respect to model size, indicating a degenerated function fitting. We conclude that the single-step approach is not as robust as the two-step approach.
Appendix D Scaling law predictions for 32B-6T
We demonstrated our method on 7B-4T and 13B-5T models, using only 1% of the compute required for the target models. Here, we test the robustness of the ladder, by trying to predict the accuracies for a much larger model of the same family: OLMo 2 32B, trained to 6T tokens (32B-6T). This model was trained with FLOPs, and the ladder models account for only 0.45% of the compute of this model.
| 32B-6T | ||||
| Task | Pred | Actual | Error | %Error |
| HellaSwag | 88.0 | 85.3 | 2.7 | 3.07% |
| ARC-Challenge | 53.6 | 66.8 | 13.2 | 24.63% |
| ARC-Easy | 77.8 | 88.8 | 11.0 | 14.14% |
| PIQA | 82.9 | 83.6 | 0.7 | 0.84% |
| CommonsenseQA | 79.5 | 75.4 | 4.1 | 5.16% |
| Social IQa | 61.1 | 62.3 | 1.2 | 1.96% |
| OpenBookQA | 45.4 | 53.2 | 7.8 | 17.18% |
| Average | – | – | 5.81 | 9.57% |
Figure 18 shows the predictions and errors for 32B-6T model. As expected, and following from B.3, the overall prediction error is higher We also observe a similar trend of higher errors for high variance tasks like ARC-C and ARC-E as with the smaller target models. However, for lower-variance tasks such as HellaSwag, PiQA, and SocialIQA, we can still predict the accuracy within an absolute error of 3 points. This indicates that the ladder models are fairly robust in their prediction abilities even as the target model is scaled to be much larger.